After robbing a bank in Dodge City, a robber gallops off at 13 mi/h. 20 minutes later, the marshall leaves to pursue the robber at 16 mi/h. how long (in hours) does it take the marshall to catch up to the robber?
Answer by checkley71(8403)

13t=16(t-1/3)
13t=16t-16/3
13t-16t=-16/3
-3t=-16/3
t=(-16/3)/-3
t=(-16/3*(-1/3)
t=16/9
t=1.7778 hours the marshall will catch the thief.
proof
13*1.7778=16(1.7778-.333)
23.11=16*1.4448
23.11=23.11